Chapter 3 Problems Saved Help Save & Exit Submit Check my work 8 You are a division manager at Toyota. If your marketing department estimates that the semiannual demand for the Highlander is -150,000-1.5P what price should you charge in order to maximize revenues from sales of the Highlander? 10 points eBook Print References Prev 8 of 9 Next>
Answer #1

ANSWER:

Toyota should charge the price which makes the demand unit elastic.

Therefore,

E (Q,P) = (-1.5) ( P / (150,000 - 1.5 P ) ) = -1

(-1.5) * ( ( P ) / ( 150,000 - 1.5 P ) ) = -1

DIVIDING BOTH SIDES BY -1.5

P / (150,000 - 1.5 P) = -1 / -1.5

P / ( 150,000 - 1.5 P) = 2 / 3

P = 2 * (150,000 - 1.5 P) / 3

P = ( 300,000 - 3 P ) / 3

P = 100,000 - P

P + P = 100,000

2 P = 100,000

P = 100,000 / 2

P = 50,000

The price that should be charged is $50,000.